# Beautiful Images, Toxic Words: Understanding and Addressing Offensive Text in Generated Images

### v2

@@ 3.2 Existing Solutions are Ineffective
Given the failure of naive solutions attempts in preventing NSFW text generation in synthetic images, we turn to existing state-of-the-art solution from the language and vision-language domain. We purely focus on methods that pursue the same goal as our work, namely making the model itself safe, such that it can be openly deployed (Suau et al., 2024; Gandikota et al., 2023; Poppi et al., 2025), rather than fixing safety issues during deployment, e.g., (Schramowski et al., 2023), which is limited to API-based models.
AURA (Suau et al., 2024). The AURA method was designed to prevent language models from generating NSFW text. Therefore, it identifies neurons that are active in toxic generation and dampens these. We adapt the method for text-to-image generation models, as detailed in Section A.2.3. In LLMs, AURA is applied to the feed-forward layers only. We perform extensive ablations to identify which layers benefit most from the intervention. Our results in Table 7 highlight that best results can be achieved when applying AURA to the text encoder’s feed-forward layers, which is in line with the original AURA method intervention and yields the insights that the text encoder might be a suitable point for our improved mitigation. To achieve the best possible results for AURA in the comparison, we report its success when performing the intervention at the text encoder’s feed-forward layers in our further experiments unless otherwise specified.
ESD (Gandikota et al., 2023). The ESD method fine-tunes the model by steering the unconditional generation away from unsafe concepts using modified classifier-free guidance while fine-tuning weights in the cross-attention and multilayer perceptron (MLP) layers. Due to its inherent reliance on a static noise schedule, it is incompatible with newer models, such as SD3 which implements a flow-matching approach (we present more details in Section A.2.4). Therefore, we assess ESD on Stable Diffusion version 1.4 (SD1.4) as done in their paper (Gandikota et al., 2023). While the inherent text-generation ability of SD1.4 underperforms SD3 significantly, applying the method still allows us to quantify the changes incurred to benign and NSFW text generation, and to assess whether ESD is an effective solution to our problem.
Safe-CLIP (Poppi et al., 2025). Safe-CLIP safety fine-tunes the CLIP model that yields the textual embeddings for DMs. It uses a custom dataset that contains unsafe images and captions with close safe counterparts and aims at mapping the unsafe inputs to their respective safe embeddings. It then fine-tunes the CLIP encoder with a combination of various losses that serve to push NSFW embeddings to a safe space, while, at the same time, preserving the embedding space on benign examples. We detail their approach further in Section A.2.5. For our experiments, we vary the weights that steer how much emphasis is put on each of the loss terms in order to assess the trade-offs between impeding NSFW generation and preserving benign performance.
@@
Experimental Setup. The full experimental setup used to implement and evaluate the baselines is presented in Section A.1. We assess the results both in terms of how the text generation changes on benign and NSFW words, and based on the quality of the generated images. A good mitigation is characterized by causing high change in the NSFW text generation (we do not want to recognize the NSFW words anymore), and a low change in the benign text generation (we want to preserve benign performance). Additionally, the overall image quality should not be significantly affected. Details on the metrics we use for evaluation are presented in Section 4.1.
Baseline Trade-offs. In Figure 3, we assess various trade-offs that can be achieved by the different baselines (for example, by applying AURA to different layers, different numbers of neurons, and with different thresholds, or by running Safe-CLIP with various weightings of the different loss terms). We find that AURA demonstrates notable inconsistency in its ability to mitigate NSFW text generation. In some setups (Aura with Dampening as detailed in Table 8) the $\Delta$1gramLD scores for both benign and NSFW are close to 0, indicating that the method fails to impact either text generation. When it does have an impact, it tends to affect both NSFW and benign text alike (the data points lie on the diagonal in Figure 3), which undermines its objective of reducing only NSFW text generation.
Figure 3: State-of-the-art Baselines. We assess the applicability of state-of-the-art baselines in mitigating the generation of NSFW text in images. Results for AURA and Safe-CLIP are obtained on SD3, ESD is applied for SD1.4 because of incompatibility with SD3. The results show that most interventions affect benign and NSFW samples proportionally, as evidenced by their alignment along the diagonal, indicating a lack of targeted toxicity mitigation, with SafeCLIP further degrading benign performance.
SafeCLIP adopts a more aggressive suppression of NSFW text, as indicated by higher $\Delta$1gramLD. However, this too comes at the cost of benign text getting affected. Additionally, SafeCLIP causes the highest image quality decrease on benign samples as reflected in its higher KID scores. On the first glance, it looks like better trade-offs are achieved using ESD. The best setup corresponds to a relatively small learning rate in the range of 1e-6 to 1e-5 during its concept removal step. Yet, when analyzing the images generated after the different interventions for NSFW and benign prompts, see Figure 7 in the Appendix, it becomes obvious that SD1.4, on which ESD is evaluated, exhibits difficulties in generating coherent text both before and after intervention, with no significant visual improvement post-intervention. Therefore, the apparent better trade-offs might be an artifact of the overall low-quality generation and are likely not to transfer to better models.
@@
When analyzing the best setup identified for each of the baseline methods in Table 2, we observe that for NSFW text, the other two methods, AURA and Safe-CLIP, exhibit an increase in NgramLev score, with AURA increasing by 2.56 and Safe-CLIP by 2.77, indicating a stronger modification of the original content. However, these modifications come at the expense of benign text generation, where AURA and Safe-CLIP also experience significant NgramLev score increase of 2.20 and 2.75, respectively. This trade-off suggests that while AURA and Safe-CLIP apply stronger transformations, they may also introduce more unintended changes to benign text. Looking at Figure 7, AURA and Safe-CLIP both still do not achieve complete removal of NSFW text, resulting in residual occurrences in the generated images. Additionally, these interventions introduce distortions in benign text generation, leading to spelling inconsistencies within the output, and indicating undesirable trade-offs.
@@
F1 LD KID CLIP-Score NgramLev F1 LD KID NgramLev
Before After $\Delta$ Before After $\Delta$ Value Before After $\Delta$ Before After $\Delta$ Before After $\Delta$ Before After $\Delta$ Value Before After $\Delta$
AURA 90 57.71 -32.29 2.30 7.70 5.40 0.062 91.70 91.48 -0.22 1.70 3.90 2.20 92.60 55.69 -36.91 1.40 10.40 9 0.063 1 1.56 2.56
ESD 33.20 28.20 -5 9.12 14.50 5.38 0.053 26.43 21.56 -4.87 3.24 5.34 2.10 31.45 27.76 -3.69 11.23 14.67 3.44 0.059 3.60 6.90 3.30
Safe-CLIP 90 65.78 -24.22 2.30 4.80 2.50 0.054 91.70 91.34 -0.36 1.70 1.05 2.75 93.00 60.12 29.88 1.40 4.68 6.08 0.058 1 1.77 2.77
Ours 90.0 60.00 -30 2.30 6.95 4.75 0.052 91.70 91.30 -0.4 1.7 2.45 0.75 92.6 69.11 22.69 1.40 5.96 4.56 0.054 1.00 3.05 2.05
Table 2: Best Baselines. We present the results for the baselines with the best parameters and our NSFW-Intervention method.
4 Our NSFW-Intervention and ToxicBench Benchmark
The shortcomings of the previous methods motivate the necessity to design methods targeted to mitigate the threat of NSFW text generation within synthetic images. To facilitate this endeavor, we introduce ToxicBench, the first benchmark to assess generative text-to-image models’ NSFW text generation ability. Finally, we propose NSFW-Intervention to prevent NSFW text generation while leaving the model’s benign and general generation abilities intact.
@@ 4.1 ToxicBench: Evaluating NSFW Text Generation
We create the ToxicBench-dataset building on CreativeBench (Yang et al., 2024b), a creative text prompt benchmark adapted from GlyphDraw (Ma et al., 2023) which holds various prompts to encourage the generation of text within the synthetic images. Examples include, ‘Little panda holding a sign that says ”<word>”.’ or ‘A photographer wears a t-shirt with the word ”<word>” printed on it.’ In total, there are 218 different prompt templates. Additionally, we curate a list of 2954 English speaking slurs, based on DirtyNaughtyList (now4real, 2025) and Toxic (Orthrus-Lexicon, 2025), two frequently updated repositories with multilingual slurs. We pre-filter the list using the roberta-base classifier from Detoxify (Hanu & Unitary team, 2020) and only keep those words that are classified as NSFW with a score above 0.9. This yields a total of 437 NSFW words. We perform a random split of the slurs into a training set with 337 and a test set with 100. This split makes sure that the same NSFW word that is seen during training does not appear at test time (with the only difference being the different prompt template). Thereby, we make sure that removing the model’s NSFW text generation ability is not only limited to the words seen during training. Then, we combine all prompt templates with the slurs, replacing the ”<word>” token. This yields a total of $73466(218\times 337)$ training data points and $21800(218\times 100)$ test data points. Finally, we provide OCR annotations with every data point for evaluation.
@@
We implement an open source pipeline to assess the models’ text generation ability and image quality. An overview of the pipeline is presented in Figure 4. Overall, our pipeline operates as follows: We start with generated images that we want to evaluate.
Figure 4: ToxicBench Evaluation Pipeline. We depict ToxicBench’s evaluation pipeline. The pipeline is designed for two main use-cases, namely 1) evaluating text and image-based metrics, for example, with the aim of assessing the impact of a mitigation method, and 2) detecting NSFW text in generated images.
Then, we perform OCR to extract all the characters embedded within. We integrate the EasyOCR model [^1] https://github.com/JaidedAI/EasyOCR, however, our pipeline can be easily extended to rely on other OCR models as well. Based on the extracted characters, there are two major use-cases: 1) Providing text and image-based metrics that can serve to assess mitigation methods. For example, if we want to assess whether the NSFW text generation ability of the model decreased through the mitigation, we have to generate the image twice, using the same prompt and random input seed, once with the model before putting the mitigation into place, and then with the protected model after mitigation. 2) The second use-case is assessing NSFW-ness of an image as standalone (as done in the right column of Table 1): In this case, after the OCR, we run a text-toxicity detector (Hanu & Unitary team, 2020) and report the score.
@@
Our metrics assess both text and overall image quality. For good mitigations, we expect that they decrease text generation for NSFW words while preserving text generation for benign words and image quality. In the following, we detail our metrics:
  • Kernel Inception Distance (KID): KID Bińkowski et al. (2018) is a metric designed to evaluate the quality of generated images by comparing their feature representations to a reference set. Unlike Fréchet Inception Distance (FID) Heusel et al. (2017), which assumes that feature embeddings follow a Gaussian distribution and relies on estimating mean and covariance matrices, KID is based on Maximum Mean Discrepancy (MMD) with a polynomial kernel. This makes KID more robust, particularly for small sample sizes, where FID can be unstable due to poor covariance estimation. KID provides an unbiased estimate of the distance between distributions, returning values in the range [0.0, 0.1], where lower values indicate higher similarity between distributions. In this setup, the reference distribution is defined by the images before intervention, and KID is computed on the images generated after intervention to quantify the impact of modifications on the generated image distribution.
[…]
  • CLIP-score: This metric for evaluating image captioning is used in our case to evaluate the overall alignment of prompt to image.
  • Levenshtein Distance (LD): The LD between the words in the text prompt and the text predicted by the OCR model measures the degree of textual similarity. Specifically, it measures the minimum number of single-character edits (i.e., insertions, deletions or substitutions) required to change one word into the other and ranges from 0 to the length of the longer string (for NSFW removal, higher is better and for benign alignment, lower is better).
  • Ngram Levenshtein Distance: We introduce this new metric building on the original LD described above. Observations have shown that generated text on images can be excessively long compared to the ground truth words from the input prompts of text-to-image models. This behavior is observed on all of the model under scrutiny. For example, when asking them to generate the word ”Newspaper”, most models also generate a sample newspaper with actual template text. In these cases, the original LD metric might not be sufficiently expressive anymore, as it would have extremely high values (due to the many insertions). Instead, we propose to replace it by our Ngram Levenshtein Distance. Our metric first divides the OCR-generated text into a several lists. Each of the lists contains all the k-adjacent tokens of the OCR-generated text. This variable k is chosen with $k\in[1,n+1]$ with $n$ the number of tokens of the ground truth text to be generated on the image. This enables to capture efficiently any substring from the original OCR-generated text which tokenization is close to the ground truth word. Finally, LD is performed on each elements of the described lists and the lowest LD is then returned.
Providing a standardized, model-agnostic evaluation with a fixed set of metrics provides a rigorous benchmark to help the community measuring NSFW text generation in images.
@@
(b) Field
(c) Laptop
(d) Radio
(e) Story
(f) Travel
@@
Figure 5: Overall NSFW-Intervention on NSFW and Benign words. Samples of generated images on the test set of ToxicBench for benign words (1st line) and NSFW words (2nd line). Overall a retaining of benign word generation is observed compared to a degradation of harmful words generation. A sample of two edge cases is presented on the right column with a spelling mistake for the word ”alley” and the highly NSFW sample ”giant cocks” is easily recognizable to the human eye.
@@ 4.2 NSFW-Intervention: Mitigating NSFW Text Generation in Images
Finally, we propose NSFW-Intervention, as an initial step toward addressing the NSFW text generation in images. Given, that the ablations on AURA suggest that the text encoder within DMs is a good points to implement the intervention (see Table 7), our NSFW-Intervention safety fine-tunes this encoder with a custom dataset. More precisely, we target the CLIP (Radford et al., 2021) encoder for its wide applicability in DMs, following (Poppi et al., 2025; Zhang et al., 2024). We find that this approach yields strong results in mitigating NSFW text while retaining the overall text and image generation ability on benign sample.
As a custom fine-tuning dataset, we extend our ToxicBench-dataset with benign replacement words for the NSFW words. We generate these words with the goal of creating a close semantic and/or grammatical correspondence to their NSFW counterparts, although being completely harmless. For instance the corresponding benign word to the NSFW word “scumbag” is “stuff bag”. The mappings are all prompted from GPT4 (OpenAI, 2024), the prompt used for obtaining the mappings is presented in Section A.1, Figure 6. Those mapped words are then integrated into our ToxicBench by replacing their corresponding NSFW words with them. Then, we instantiate a loss function with the objective of mapping NSFW prompts to their benign counterparts. We refer to the NSFW prompts as $x_{NSFW}$ and their generated counterparts as $x_{benign}$. We define this loss as:
$\text{NSFWLoss}(x_{NSFW},x_{benign})=$ (1)
$\text{CosSimLoss}(\hat{M}(x_{NSFW}),M^{*}(x_{benign}))$
with $\hat{M}$ and $M^{*}$ the fine-tuned and the frozen CLIP text encoders respectively.
Our choice of safety fine-tuning setup through mapping NSFW words to semantically close words has two major motivations: 1) A negative loss setup for “forgetting” NFSW text embeddings in the CLIP Embedding space is hard to implement. We experimented with this setup and found that this loss is very small ($\approx 10^{-9}$) which makes it too small for training because of computational precision. In contrast, the loss defined by Equation 1 is large enough to be computed without instability caused by this computations approximations. 2) The semantic closeness between the NSFW word and its replacement aims at lowering the initial training loss by using the properties of the CLIP embedding space. This makes learning succeed faster, and is better for mitigating utility drop on the benign samples.
Our NSFW-Intervention Outperforms the Baselines.
We empirically evaluate our NSFW-Intervention using our ToxicBench benchmark. The full experimental setup is specified in Section A.1.
We solely apply NSFW-Intervention on SD3, due to its reliance on CLIP encoders and its strong text generation ability. Since SD3 leverages two CLIP text encoders $E_{1}$ and $E_{2}$, we need to safety fine-tune both of them simultaneously. To reduce additional sources of potential NSFW behavior, we do not include the T5 model used by the original SD3, which was reported by Esser et al. (2024) to introduce marginal improvements in the model’s generation ability. The best fine-tuning hyperparameters for both encoders, identified through grid-search, are specified in Table 3. Once, safety fine-tuning terminates, we evaluate the success of our NSFW-Intervention. The empirical results in Figure 3 highlight that NSFW-Intervention outperforms the baselines by decreasing the NSFW text generation ability more than twice as much as the benign text generation ability, and thereby, being above the diagonal in the plot. Additionally, when looking into the images generated after our NSFW-Intervention, in Figure 5, and comparing them to the samples generated after applying the baseline methods, see Figure 7, we observe that NSFW-Intervention is able to maintain benign text generation, and to mitigate NSFW text generation. We would like to note again that, due to the strict train and test split in our data, as described in Section 4.1, none of the NSFW words from the prompts in Figure 5 were seen during training. This highlights our approach’s ability to mitigate text-to-image models’ general NSFW text generation.

### v3

@@ 3.2 Existing Solutions are Ineffective
Given the failure of naive solution attempts in preventing NSFW text generation in synthetic images, we turn to existing state-of-the-art solutions from the language and vision-language domains. We purely focus on methods that pursue the same goal as our work, namely making the model itself safe, such that it can be openly deployed (Suau et al., 2024; Gandikota et al., 2023; Poppi et al., 2025), rather than ensuring safety during deployment (Schramowski et al., 2023), which is limited to API-based models.
AURA (Suau et al., 2024). We adapt the AURA method, originally developed to suppress toxic generation in LLMs by dampening neurons in feed-forward layers, to DMs (see Section A.11.3). Through ablations presented in Table 16, we find that applying AURA to the text encoder’s feed-forward layers yields the best results, consistent with the original method. Unless stated otherwise, all experiments apply AURA at this location.
ESD (Gandikota et al., 2023). ESD fine-tunes DMs by steering unconditional generation away from unsafe concepts using a modified classifier-free guidance loss, targeting cross-attention and MLP layers. Since ESD relies on a fixed noise schedule, it is incompatible with SD3’s flow-matching framework. As in the original paper, we evaluate ESD on SD1.4 and report its effect on NSFW and benign text generation. Implementation details are in Section A.11.4.
Safe-CLIP (Poppi et al., 2025). Safe-CLIP fine-tunes a CLIP encoder to push unsafe prompts toward safe embedding regions using contrastive losses over paired NSFW and benign prompts. We adopt their setup as described in Section A.11.5, which includes implementation details and dataset construction. We sweep loss weights to assess trade-offs between NSFW suppression and benign preservation, and report results using the best-performing configuration.
@@
Setup and Evaluation. The full experimental setup used to implement and evaluate the baselines is presented in Section A.11. We assess the results both in terms of how the text generation changes on benign and NSFW words, and based on the quality of the generated images. A good mitigation is characterized by causing high change in the NSFW text generation (we do not want to recognize the NSFW words anymore), and a low change in the benign text generation (we want to preserve benign performance). We measure these changes in the number of deleted, added, and substituted characters after the intervention with a new dedicated metric we propose, namely the N-gram Levenstein Distance (NGramLD). A good mitigation achieves low NGramLD for benign words and high NGramLD for NSFW words, indicating few or many changes to the words, respectively. Details on the metrics we use for evaluation are presented in Section 4.1. Finally, we require a good mitigation to not affect the overall image quality significantly.
@@
Baseline Trade-offs. When analyzing the best setup identified for all of the baseline methods in Table 2, we observe that for NSFW text, AURA and Safe-CLIP cause an increase in NGramLD score. AURA increases the score by 2.56 and Safe-CLIP by 2.87. This suggest that both are effective in making the NSFW words less recognizable, as we also show visually in Figure 10. However, these modifications come at the expense of benign text generation, where AURA and Safe-CLIP also experience significant NGramLD score increase of 2.20 and 2.65, respectively, i.e., the methods affect the benign text nearly as much as the NSFW text. This suggest that they cause more of an overall text quality degradation rather than a targeted NSFW text quality mitigation. More extensive results for applying AURA to the other evaluated DMs can be found in Section A.11.3. Compared to AURA and Safe-CLIP, we observe the best baseline trade-off with ESD on SD 1.4, with NGramLD increasing of only 2.10 for benign text and 3.30 for NSFW text. But, as demonstrated by the very high values of Levenshtein Distance (LD) for benign and NSFW text generation (14.50 and 14.67 respectively) and the low CLIP-Score compared to other baseline methods, the overall image quality of SD 1.4 is very low, diminishing the relevance of the results to our present study. More details about the limits of the applicability of ESD to text mitigation in images are presented in Section A.11.4. Overall, Figure 10 suggests that neither of the baselines achieves complete removal of NSFW text. Additionally, they introduce distortions in benign text generation, leading to spelling inconsistencies within the output, and indicating undesirable trade-offs.
@@
LD KID CLIP-Score NGramLD LD KID NGramLD
Before After $\Delta\downarrow$ Value Before After $\Delta$ Before After $\Delta\downarrow$ Before After $\Delta\uparrow$ Value Before After $\Delta\uparrow$
ESD 9.12 14.50 5.38 0.053 26.43 21.56 -4.87 3.24 5.34 2.10 11.23 14.67 3.44 0.059 3.60 6.90 3.30
AURA 2.30 7.70 5.40 0.062 91.70 91.48 -0.22 1.70 3.90 2.20 1.40 10.40 9.00 0.063 1.00 1.56 2.56
Safe-CLIP 2.30 8.90 6.60 0.068 91.70 87.43 -4.27 1.70 0.95 2.65 1.40 9.34 7.94 0.063 1.00 1.87 2.87
Table 2: Best Baselines. We present the results for the baselines with the best parameters. Up and down arrows indicate the preferred (higher or lower) changes in evaluation metrics after intervention.
4 Our ToxicBench Benchmark and NSFW-Intervention
The shortcomings of the previous methods motivate the necessity to design methods targeted to mitigate the threat of NSFW text generation within synthetic images. To facilitate this endeavor, we introduce ToxicBench, the first benchmark to assess generative text-to-image models’ NSFW text generation ability. Additionally, we propose NSFW-Intervention to prevent NSFW text generation while leaving the model’s benign and general generation abilities intact.
@@ 4.1 ToxicBench: Evaluating NSFW Text Generation
We create the ToxicBench dataset consisting of 218 prompt templates adapted from CreativeBench (Yang et al., 2024b) each designed to elicit visible text in generated images (e.g., ‘Little panda holding a sign that says "<word>".’). We curate 437 NSFW words using Detoxify (Hanu and Unitary team, 2020), and pair each with a benign alternative generated by GPT-4 that is semantically close. These are split into 337 training and 100 held-out test pairs to evaluate generalization on unseen NSFW words. Combined with the prompt templates, this yields 73466 training and 21800 test prompt pairs. We refer to Section A.5 for a comprehensive description of ToxicBench.
@@
We implement an open source evaluation pipeline to assess both the textual content and visual quality of generated images. An overview is shown in Figure 3. The pipeline begins with a generated image and applies OCR using EasyOCR [^1] https://github.com/JaidedAI/EasyOCR to extract any visible text. The pipeline is modular and can be extended to alternative OCR models. Based on the extracted text, the pipeline supports two use cases:1) Mitigation Evaluation: We generate two images using the same prompt and random seed: one before and one after applying the mitigation. This allows us to directly compare the changes in embedded text and image quality using our evaluation metrics. 2) Standalone Detection: We evaluate a single image by running a toxicity classifier (Hanu and Unitary team, 2020) on the OCR output to determine whether it contains harmful text (e.g., as in the right column of Table 1).
@@
Our evaluation metrics assess both the quality of generated images and the fidelity of rendered text. Effective mitigation should reduce the presence of NSFW text without degrading the image quality or suppressing benign content. We use the following metrics:
Figure 3: ToxicBench Evaluation Pipeline. The pipeline is designed for two main use-cases, namely 1) evaluating text and image-based metrics, for example, with the aim of assessing the impact of a mitigation method, and 2) detecting NSFW text in generated images.
  • Kernel Inception Distance (KID) and CLIP-Score: Our image quality evaluation metrics. KID (Bińkowski et al., 2018) measures the distributional distance between generated images (after intervention) and a reference set (before intervention) based on features extracted from an Inception network. CLIP-Score evaluates the overall alignment between a given prompt and image.
  • Levenshtein Distance (LD): LD measures the minimum number of single character edits (insertions, deletion, or subtitutions) required to transform the target word into the OCR-extracted text. For NSFW prompts, a higher LD is desired (indicating disruption of NSFW text); for benign prompts a lower LD reflects preservation of the intended word.
  • Ngram Levenshtein Distance (NGramLD): Given that DMs often embed long sequences (e.g., generating full newspaper layouts when prompted with ’Newspaper’) in the generated image, standard LD can be overly penalizing. Therefore, we introduce a modified version of LD, namely NGramLD. Our new metric first extracts all $k$-grams ($k\in[1,n+1]$, where $n$ is the number of tokens in the ground truth word) from the OCR output. We then compute LD between the ground truth word and each $k$-gram substring, returning the minimum score. This method robustly detects partial matches while avoiding bias toward long OCR strings, since it compares only the most relevant substrings rather than penalizing the full text length.
@@
(b) Cub
(c) Belt
(d) Hill
(e) Shark
(f) Truck
@@
Figure 4: Overall NSFW-Intervention on NSFW and Benign words. Samples of generated images from SD3 on the test set of ToxicBench for benign words (1st line) and NSFW words (2nd line). We present 2 edge cases on the right column with a spelling mistake for the word "puzzle" and the highly NSFW sample "giant cocks" is easily recognizable to the human eye.
@@ 4.2 NSFW-Intervention: Mitigating NSFW Text Generation in Images
Next, we introduce NSFW-Intervention, our novel and generalizable method for mitigating NSFW text generation in images. NSFW-Intervention directly fine-tunes the backbone of DMs to alter the visual representation of NSFW language. It leverages supervision from ToxicBench to perform targeted intervention—modifying only the rendering of harmful words while preserving overall image quality and text generation for benign inputs. Using only NSFW words from ToxicBench, NSFW-Intervention generalizes its mitigation ability to all tested NSFW words.
1. A Carefully Curated Fine-Tuning Dataset.
To train a model that avoids generating NSFW text while preserving the rest of the image, we construct a fine-tuning dataset specifically for this goal. Starting from ToxicBench, we use NSFW prompts to generate images that contain harmful embedded text. For each prompt, we then replace the NSFW word with a carefully chosen benign counterpart and regenerate the image using the image editing method of Staniszewski et al. (2025). This involves caching intermediate activations from the first (NSFW) generation and reusing them during generation with the benign prompt, resulting in nearly identical image pairs that differ only in the rendered text. We collect these samples into training triplets $(x_{\text{NSFW}},I_{\text{NSFW}},I_{\text{benign}})$, where $x_{\text{NSFW}}$ is the original NSFW prompt, and $I_{\text{NSFW}},I_{\text{benign}}$ are the two images that differ only in embedded text. This dataset serves our training objective: generate the same image structure from a NSFW prompt, but with benign text instead of harmful content.
2. A Targeted Safety Fine-Tuning Approach.
For fine-tuning with the dataset described above, we build on recent findings by Staniszewski et al. (2025), which show that text rendering in DMs is localized to a small subset of attention layers. By restricting updates to only these layers, we can suppress harmful text while preserving general image generation quality. This targeted strategy also reduces the number of trainable parameters and minimizes interference with unrelated visual content. Notably, our method fine-tunes the generative backbone rather than the text encoder. [^2] We also experimented with fine-tuning the CLIP encoder present in some DMs. However, this approach is not applicable across all architectures and consistently underperformed our backbone-level intervention. See Section A.7 for details.
At each training step, we start with an image $I_{\text{NSFW}}$ containing harmful embedded text, generated from an NSFW prompt. This image is corrupted with Gaussian noise at a randomly sampled diffusion timestep $t$, where larger $t$ values correspond to noisier images and $t=0$ to the fully denoised one, yielding $I_{\text{NSFW}}(t)$. The model is tasked with predicting the denoised output, conditioned on the NSFW prompt embedding $\phi(x_{\text{NSFW}})$, but is trained to match a benign target image $I_{\text{benign}}$ that retains the same visual structure but replaces the harmful text. By training on a diverse set of NSFW prompts and their safe counterparts, the model learns to suppress a broad range of harmful text patterns, including those not seen during training.
To guide this training more effectively, we vary the denoising timestep $t$, ensuring that suppression is learned progressively throughout the generation process. This allows the model to influence the emergence of harmful tokens even in early stages. To emphasize correction when the text is most visible, we apply the standard timestep-dependent weight $w(t)$ that increases as $t$ approaches 0. In our implementation, $w(t)$ follows a logit-normal schedule: timesteps are normalized to $[0,1]$, passed through a logit transformation, and evaluated under a normal distribution with mean $\mu=0$ and standard deviation $\sigma=1$. This yields a weighting curve that prioritizes mid-to-late denoising steps, where embedded text becomes clearest. The full training objective is
$\mathcal{L}(x_{\text{NSFW}},I_{\text{NSFW}}(t),I_{\text{benign}},t)=\left\|w(t)\cdot\left(f_{\theta}(I_{\text{NSFW}}(t),t,\phi(x_{\text{NSFW}}))-I_{\text{benign}}\right)\right\|^{2}\text{,}$ (1)
where:
  • $\phi(x_{\text{NSFW}})$ is the frozen text encoder’s embedding of the NSFW prompt,
  • $w(t)$ is a timestep-dependent weight emphasizing denoising steps close to $t=0$,
  • $f_{\theta}(I_{\text{NSFW}}(t),t,\phi(x_{\text{NSFW}}))$ is the predicted denoised image after one step.
This loss encourages the model to align its denoised prediction with the benign target image, despite being conditioned on the original NSFW prompt. At inference, this enables suppression of harmful text while preserving the surrounding visual content.
Note that while NSFW-Intervention is designed for DMs, it can also be easily extended to the novel Visual Autoregressive Models (VARs) (Tian et al., 2024; Tang et al., 2024). We show this extension and an evaluation on the state-of-the-art Infinity (Han et al., 2024) model in Section A.10.
5 Results
Benign Text NSFW Text
LD KID CLIP-Score NGramLD LD KID NGramLD
Before After $\Delta\downarrow$ Value Before After $\Delta$ Before After $\Delta\downarrow$ Before After $\Delta\uparrow$ Value Before After $\Delta\uparrow$
SD3 2.30 7.92 5.62 0.059 91.70 72.90 -18.80 1.70 6.50 4.80 1.40 8.37 6.97 0.061 1.00 7.90 6.90
DeepFloyd IF 3.76 9.56 5.80 0.059 90.98 75.40 -15.58 1.82 5.36 3.54 2.70 9.56 6.86 0.060 1.89 6.67 4.78
SDXL 5.67 11.78 6.11 0.063 88.72 68.71 -20.01 2.37 7.00 4.63 5.90 14.65 8.75 0.065 2.14 7.85 5.71
Table 3: NSFW-Intervention. We present the results for NSFW-Intervention.
NSFW-Intervention Mitigates NSFW Text While Preserving Image Quality.
We evaluate our method on the ToxicBench benchmark across multiple DMs (the detailed experimental setup can be found in Section A.4) . As shown in Table 3, our method improves the trade-off between suppressing NSFW text and preserving benign outputs across all models. For example, on SD3, our method increases the suppression of NSFW text, improving the NGramLD from 1.18 with AURA to 6.90, while preserving benign generation with a score of 4.80, leading to a +2.10 differential NGramLD value between NSFW/benign text. Similar improvements are observed on DeepFloyd IF and SDXL, where harmful content is more effectively suppressed (+1.09 and +1.08 respectively) without sacrificing benign quality. Despite strong mitigation, NSFW-Intervention maintains image quality: the KID score increases by at most 9% across benign samples, and FID scores show minimal degradation (Section A.9). Qualitative results (Figure 4) illustrate that NSFW terms are rendered unreadable while benign text remains legible, with similar trends in SDXL and DeepFloyd IF (Figure 5 and Figure 6).
Prompt Type Before Intervention After Intervention
NSFW 78.67±1.12 26.56±1.07
Misspelled NSFW 76.41±1.12 11.45±1.02
Benign 83.43±1.15 55.40±1.04
Table 4: User Study. The table reports mean accuracy with standard deviation (mean±$\pm$±std). Our intervention significantly reduces perceived toxicity for NSFW prompts, with a moderate effect on benign prompts. Results on misspelled NSFW prompts fall between the two, demonstrating robustness even against character-level obfuscation.
Ablation Study.
To assess the importance of layer selection in effective mitigation, we ablate the design by applying NSFW-Intervention uniformly across all joint (SD3) and cross-attention layers (SDXL, DeepFloyd IF), rather than restricting updates to those used in text generation. As detailed in Table 11, this broader intervention results in substantially weaker suppression of NSFW text. On SD3, for instance, NGramLD improves by only +0.49, compared to +6.90 when updates are limited to the text-generation layers (Table 3). Similar trends are observed on SDXL and DeepFloyd IF. This highlights that targeting specific layers yields substantially stronger mitigation results.
A user study demonstrates the effectiveness of NSFW-Intervention.
To evaluate the impact of our intervention, we conducted a user study measuring how participants perceived generated text before and after its application (see Section A.6 for details). Participants labeled images from NSFW, benign, and misspelled NSFW prompts as either safe or unsafe, and rated benign text as readable or unreadable. As shown in Table 4, recognition accuracy for NSFW prompts dropped from 78.67% to 26.56% after intervention, indicating a substantial reduction in the readability of harmful text. The effect was even stronger for misspelled NSFW prompts, which were not included during training; accuracy dropped from 76.41% to just 11.45%, highlighting strong generalization to adversarial variants. Meanwhile, benign text remained relatively readable, with post-intervention recognition at 55.40%, more than twice that of NSFW and nearly five times that of misspelled NSFW prompts. These findings demonstrate NSFW-Intervention’s ability to suppress harmful outputs while preserving benign content, even under distributional shifts.